OpenStudy (anonymous):

- tan^2x + sec^2x = 1

OpenStudy (anonymous):

you need the equation for the tan and sec to be able to solve this

OpenStudy (anonymous):

I don't understand how to use the equations or identities or whatever

zepdrix (zepdrix):

Pythagorean Identity for Tangent: \(\large\rm \color{orangered}{\tan^2x+1=\sec^2x}\) Do you see how you can make use of that in your problem? :)\[\large\rm - \tan^2x + \color{orangered}{\sec^2x} = 1\]

OpenStudy (anonymous):

No, I'm really confused.

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

This is your Pythagorean Identity,\[\large\rm \color{orangered}{\tan^2x+1=\sec^2x}\]You'll just have to acecpt that it's an identity for now. Notice, it tells us that \(\large\rm \sec^2x\) is the same thing as \(\large\rm \tan^2x+1\) So in our problem here,\[\large\rm - \tan^2x + \color{orangered}{\sec^2x} = 1\]we can make the exchange,\[\large\rm - \tan^2x + \color{orangered}{\tan^2x+1} = 1\]

zepdrix (zepdrix):

And then simplify.

OpenStudy (anonymous):

So 1=1 since -tan^2x + tan^2x cancels out, right

zepdrix (zepdrix):

yay good job \c:/

zepdrix (zepdrix):

What's really important in this type of problem, is that you `only manipulate one side of the equation`. It's not like normal algebraic stuff where we add, subtract across the equal sign. We have to leave one side alone completely. We didn't mess with the 1 on the right in this problem, so we did good!

OpenStudy (anonymous):

Thanks very much!